Circular No. 6708

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cz IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (V = 20.5) on V and R CCD images taken on July 19 by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cz is located near Abell 3565 at α = 13^{h}37^{m}09^{s}.39, δ = −33^{o}47′31′′.9 (equinox 2000.0). A nearby star has position end figures 13^{s}.08, 36′′.4. SN 1997cz was also present at approximately the same magnitude on similar V and R frames taken on July 21 and June 10, but it is not visible on images taken on Apr. 26 (limiting mag R = 21).

SUPERNOVA 1997cy IN ANONYMOUS GALAXY S. Benetti and A. Pizzella, European Southern Observatory; and P. J.

Wheatley, University of Leicester, report: “A fully-reduced CCD spectrum (range 320–880 nm, resolution 2.0 nm) of SN 1997cy (cf. IAUC 6706), obtained on July 24.43 UT with the Danish 1.5-m telescope (+ DFOSC), is dominated by an intense emission line with a composite profile. Its narrow component is centered at 698.4 nm (FWHM = 4.0 nm), and it is superimposed on a broader one (centered at ∼ 697.0 nm, FWHM ∼ 13.0 nm) that has a very steep red wing. If this feature is identified with Hα, then a redshift of 0.0642 is deduced for this supernova. The bluer part of the spectrum shows strong and broad Fe II features resembling those shown by some type-Ic supernovae near maximum; peaks are measured at ∼ 493.0 nm (FWHM ∼ 15.8 nm), 529.0 nm (FWHM ∼ 12.0 nm), and 580.0 nm (FWHM ∼ 45.0 nm). Another very broad emission (FWHM ∼ 70.0 nm) is centered at ∼ 785.5 nm. A relatively faint Hβ emission is also present (peak at 518.0 nm, FWHM = 2.7 nm). Spectroscopic and photometric follow-up of this peculiar type-II supernova is encouraged.”

BL LACERTAE F. Makino, J. Mattox, T. Takahashi, and J. Kataoka, Insitute of Space

and Astronautical Science (ISAS); and H. Kubo, Physical and Chemical Research Institute (Riken), Wako, report: “ASCA observed this source during July 18.60–19.62 UT, yielding an average x-ray flux of 2.6 × 10^{−11} erg s^{−1} cm^{−2} (2–10 keV), which is more than four times higher than that observed in its low state with Ginga (Kawai et al. 1991, Ap.J. 382, 508). The spectrum was of power-law type with photon index of −1.44 ± 0.01. The flux showed variation by a factor of 2 during the observation.”

1997 July 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green